Grade 11General Physics

Two vectors a and b are at an angle of 60 degree
with each other. Their resultant makes an angle of 45degree with a. If b=2units, then magnitude of vector a is

Tvsvfsvfsfvrs
8 Years agoGrade 11

2 Answers

Arun
8 Years ago
Dear student
 
As we have learnt a formula for resultant angle
tan @  = a sin theta / b + a cos theta
tan 45 = a sin 60 /( 2 +  a cos 60)
2 + a/2 = sqrt (3) * a/2
a(sqrt3 - 1) /2 = 2
a = 4 / (sqrt(3) - 1)
ankit singh
5 Years ago

The vectors make 60 degrees with each other, and the resultant vector  makes 45 degrees with a

 

The other angle = 15 degrees

 

Using the Sine Rule:

 

a/sin15 = 2/sin45

 

a = 2Sin15/Sin45

 

a = 0.732